OpenStudy (anonymous):

Alice purchased paint in a bucket with a radius of 9 in. and a height of 9.5 in. The paint cost $0.09 per in3. What was the total cost of the paint, rounded to the nearest cent? Use 3.14 to approximate pi. $

OpenStudy (anonymous):

@Hoslos @whydoihavetosignup1 @Michele_Laino @abb0t @mikkel900 @bibby @perl @thomaster @nincompoop @ayyookyndall @sweetburger

OpenStudy (michele_laino):

here we have to compute the volume V of our bucket. Now a volume V of a cylinder is giben by the subsequent formula: \[V = \pi \times {r^2} \times h\] where r is the radius of our buck, and h is the heoght of our buck. so what is the volume V?

OpenStudy (anonymous):

The volume would be 2416.23 @Michele_Laino

OpenStudy (anonymous):

correct? @Michele_Laino

OpenStudy (michele_laino):

ok! that's correct. Now we have to multiply that volume by 0.09, namely: 2416.23*0.09=...?

OpenStudy (anonymous):

217.46

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

THanks :)

OpenStudy (michele_laino):

thanks! :)